# Deblur-Avatar: Animatable Avatars from Motion-Blurred Monocular Videos

### v1

@@ Abstract
We introduce Deblur-Avatar, a novel framework for modeling high-fidelity, animatable 3D human avatars from motion-blurred monocular video inputs. Motion blur is prevalent in real-world dynamic video capture, especially due to human movements in 3D human avatar modeling. Existing methods either (1) assume sharp image inputs, failing to address the detail loss introduced by motion blur, or (2) mainly consider blur by camera movements, neglecting the human motion blur which is more common in animatable avatars. Our proposed approach integrates a human movement-based motion blur model into 3D Gaussian Splatting (3DGS). By explicitly modeling human motion trajectories during exposure time, we jointly optimize the trajectories and 3D Gaussians to reconstruct sharp, high-quality human avatars. We employ a pose-dependent fusion mechanism to distinguish moving body regions, optimizing both blurred and sharp areas effectively. Extensive experiments on synthetic and real-world datasets demonstrate that Deblur-Avatar significantly outperforms existing methods in rendering quality and quantitative metrics, producing sharp avatar reconstructions and enabling real-time rendering under challenging motion blur conditions.
@@ 1 Introduction
Current human avatar modeling approaches neglect motion blur, which can severely degrade the quality of reconstructed avatars by obscuring dynamic human motion details. Several works have proposed to reconstruct static scenes with camera motion-blurred multi-view inputs, restoring a sharp NeRF or 3DGS by deformable kernels [46, 32, 58] and camera motion model [81, 97]. A dynamic pipeline [75] aims to reconstruct dynamic scenes from blurred monocular videos, but its deblurring paradigm is the same as the previous pipelines [81, 97]. These methods perform well on their datasets, however, these methods, including the dynamic one [75] are primarily designed to handle camera-induced motion blur. In their deblurring scenarios, the captured image is blurred in the pattern of large camera movements, and the objects of the scene have limited or no movements.
In contrast, in human avatar modeling tasks, motion blur is primarily caused by human movements instead of camera motions. To address motion blur in monocular dynamic human videos, we aim to reconstruct sharp 3D human avatars by explicitly modeling human movements. We propose Deblur-Avatar, an animatable avatar framework to incorporate a human movement-based motion blur model into 3DGS. We integrate human motion dynamics into 3DGS training by utilizing an interpolation function to model virtual human motion trajectories over the camera’s exposure period. During training, these motion trajectories are optimized jointly with the Gaussians, enabling the reconstruction of sharp human avatars despite motion blur. Specifically, the trajectory of each motion-blurred image is represented by the initial and final human body poses at the beginning and end of the exposure time, respectively. Given that the exposure time is relatively short, we interpolate between these initial and final poses to determine each intermediate pose along the trajectory. Using this trajectory, we generate a sequence of virtual sharp images by deforming the Gaussians with non-rigid and rigid transforms and projecting the Gaussians onto the image plane. These virtual sharp images are then averaged to synthesize the blurred image, adhering to the physical principles of motion blur. Finally, we observe that human movement does not involve the whole body for each frame. Therefore, we propose pose-dependent fusion to effectively distinguish the blurred regions for each pose in a frame, so that our model can focus on modeling motion blur in these blurred areas.
To address the absence of existing human avatar motion deblurring datasets, we collect synthetic and real-world datasets and conduct an extensive analysis to demonstrate that Deblur-Avatar effectively addresses motion-blurred inputs and renders high-quality sharp avatars, outperforming existing methods. Additionally, we conduct ablation studies to validate the effectiveness of each component. In summary, our main contributions are as follows.
@@
    We present the first framework designed to recover sharp animatable human avatars from monocular videos with motion blur. Unlike previous deblurring pipelines, we are the first to focus on motion blur from human movements.
@@ 2 Related Work
A disadvantage of NeRF-based methods is their inefficiency in training and rendering. To reduce computational costs, multi-resolution hash encoding [11, 22], empty-space skipping [22], point-based rendering [65, 90, 99, 98], UV Volumes [5] as well as dynamic MLP maps [61] have emerged as efficient alternatives. 3D Gaussian Splatting [24] (3DGS) shows superiority in quality and efficiency, so 3DGS is integrated into human avatars [102, 15, 20, 26, 35, 39, 44, 43, 50, 77]. KL dispersion [16] and a segmentation plus scaling strategy [36] are used to manipulate point cloud, and kinematic information [80] is used for surface deformation. Non-rigid deformations [62, 84] are integrated into 3DGS to achieve a balance between quality and speed. Lifted optimization and a novel differentiable shading are introduced  [84, 70] to optimize Gaussian parameters while projecting them onto meshes. Contrastive learning [95] and semantic priors [96] are used to capture details and maintain consistency.
@@
Motion Deblurring. Traditional single-image deblurring models the blur with normalized gradient sparsity [28], unified probabilistic model [69], and rotational velocity [86]. Several methods [51, 29, 78, 30] train neural networks on datasets [71], exploring attention mechanisms [92, 79] and multi-scale pipelines [64, 9]. On the other hand, video deblurring harnesses temporal information among different frames to ensure consistency. Conventional approaches employ optical flow to model inter-frame relationships [19, 55]. Several approaches stack neighboring frames as the input of neural networks [73, 100]. Attention mechanisms [94], motion magnitude [83], and wavelet-based [56] methods are proposed to improve performance and efficiency. While these methods yield excellent results, they assume the blur is predominantly from camera movement, and the method performs poorly on blur from human motion. Furthermore, these methods operate in image space and fail to capture scene geometry in $3$D space.
@@
Recently NeRF has demonstrated impressive results in novel view synthesis, and many methods have extended NeRF from sharp static scenes to motion-blurred static [46, 32, 81, 33] and dynamic scenes [75]. These methods restore sharp radiance fields by modeling motion blur from blur kernels  [46, 32] or camera motions [81, 75]. To improve efficiency, 3DGS is also introduced to model the blur with the covariance of 3D Gaussian [31, 7, 58] and camera motions [3, 97, 34]. Current methods, including the dynamic deblurring pipeline, are designed specifically for motion blur from camera motions instead of human motions, and would not suffice for our aim: restoring from motion blur due to human movements in human avatar modeling.
3 Method
@@
Figure 2: Pipeline of Deblur-Avatar. For each frame with its corresponding body pose as input, we introduce human motion trajectory modeling to predict human poses $\theta_{start}$ and $\theta_{end}$ at the start $T_{start}$ and end $T_{end}$ of the exposure period of the frame. Then we use interpolation to obtain a virtual human pose sequence. With the pose of each virtual timestamp, we perform non-rigid deformation and rigid transformation to the 3D Gaussian initialized in canonical space. The observation space virtual 3D Gaussians and their colors are accumulated through differentiable rasterization to render sharp virtual images. These images are averaged as Eq. 6 to produce the blurred result. The blurred image is then blended with the virtual sharp result by pose-dependent fusion to render the final output, supervised by the captured input RGB image. During testing, Gaussians can be rendered directly to produce a sharp image without motion trajectory modeling and pose-dependent fusion, enabling fast rendering.
@@
We propose Deblur-Avatar, a human avatar modeling pipeline to recover sharp novel views and poses from a motion-blurred monocular video. In the following, we first revisit the concept of 3DGS and the principle of motion blur. In section 3.3, we propose human motion trajectory modeling to account for human movements in the exposure period. We show how to perform rigid and non-rigid deformation from the canonical space in section 3.4. Then we apply pose-dependent fusion to predict the blurred region of each view, blending the rendered blur region and the rendered sharp region to get optimal results in section 3.5.
3.1 Preliminary: 3D Gaussian Splatting
@@
3.2 Preliminary: Motion Blur Formation
@@
3.3 Human Motion Trajectory Modeling
@@
We aim to estimate both $\theta_{\text{start}}$ and $\theta_{\text{end}}$ for each frame, along with the learnable parameters of Gaussians $G$. To predict both poses from a single input, we initialize $\theta_{\text{start}}$ and $\theta_{\text{end}}$ with the input body pose parameters, then we add random perturbation on these two sets of poses. The perturbed poses are used for subsequent rigid transformation, allowing direct optimization through backpropagation. For very small angles (i.e., when $\sin(\phi)$ is close to zero), we assume the corresponding frame has very little motion blur and linear interpolation is used as an approximation. We also predict the sequences of per-frame translations with direct linear interpolation and apply Slerp to the global orientation of the body to get a virtual sequence. We set the virtual frame number $n=5$ as the number of images during an exposure time, balancing training efficiency and rendering quality.
3.4 Human Deformation and Color Prediction
@@
3.5 Pose-dependent Fusion
@@
where this mask is used to blend the virtual sharp image $\mathbf{C}(x,j)$ with the blurred image $\mathbf{B}(x,j)$ as follows:
@@
3.6 Optimization
We learn the Gaussians and the human motion trajectories for each frame from the loss functions, consisting of an RGB loss $L_{\text{rgb}}$ and perceptual loss $L_{\text{percept}}$ for $\mathbf{C}_{\text{out}}(x,j)$ and $\mathbf{B}_{\text{gt}}$, a mask loss $L_{\text{mask}}$ for foreground human. Skinning weight regularization $L_{\text{skin}}$ and isometric regularization $L_{\text{isopos}},L_{\text{isocov}}$ [62] are implemented to alleviate noisy deformation. The full loss function is defined as:
$\begin{split}L_{\text{full}}&=L_{\text{rgb}}+\lambda_{1}L_{\text{percept}}+\lambda_{2}L_{\text{mask}}+\lambda_{3}L_{\text{skin}}\\ &+\lambda_{4}L_{\text{isopos}}+\lambda_{5}L_{\text{isocov}}\,.\end{split}$ (18)
4 Experiments
@@
Figure 3: Qualitative results on ZJU-MoCap-Blur. Compared with existing human avatar methods, our method generates sharper novel views and preserves more details with original motion-blurred inputs. Refer to supplementary for results on video deblurring preprocessing.
Figure 4: Qualitative results on Real-Human-Blur dataset. Our method generates sharper novel poses that are more faithful and preserve more details. We show more results in supplementary.
Table 1: Quantitative results on ZJU-MoCap-Blur. The input sequence is the motion-blurred images. The best performance is in boldface, and the second best is underlined. We present the training time as GPU and the rendering speed as FPS.
@@
Table 2: Quantitative results on ZJU-MoCap-Blur (pre-deblurred). The input sequences are deblurred by state-of-the-art video deblurring method [56]. The best performance is in boldface, and the second best is underlined. We present the training time as GPU and the rendering speed as FPS.
@@ 4.1 Datasets
ZJU-MoCap-Blur. This is the main dataset for quantitative evaluation. We pick six sequences (377, 386, 387, 392, 393, 394) from the ZJU-MoCap dataset and follow the training/test split of HumanNeRF [85]. We synthesize motion blur as the pipeline of current benchmark datasets on motion deblurring [51, 101, 52]. We choose state-of-the-art frame interpolation method [93] to create a high frame rate, then we average this sharp high frame rate successive frames to generate a blurry image to approximate a long exposure time, and the generated frames are temporally centered on a real-captured ground truth frame from original ZJU-MoCap. The number of averaging frames (17,33,49) is randomly assigned for each sequence, and the varying numbers indicate different blur sizes. To make the dataset more realistic, we use EasyMocap [8, 59] to re-calculate the human poses and the human masks of the synthesized blurred image sequences. Following previous works [62, 85], we conduct quantitative evaluations on novel view synthesis and show qualitative results for animation on out-of-distribution poses. LPIPS in all the tables are scaled up by 1000.
@@
Real-Human-Blur. Due to the lack of real-world datasets specifically designed for addressing motion blur in human avatar modeling, we have curated a dataset consisting of monocular motion-blurred videos. This dataset was captured using a DSLR camera, along with an additional video sourced from the internet. We use SPIN [27] to obtain approximate body poses and employ SAM [63] for segmenting the foreground human. Since there is no accurate sharp ground truth for these real captures, we use this dataset solely for qualitative comparisons on novel pose synthesis.
4.2 Baseline Comparisons

### v3

@@ Abstract
We introduce a novel framework for modeling high-fidelity, animatable 3D human avatars from motion-blurred monocular video inputs. Motion blur is prevalent in real-world dynamic video capture, especially due to human movements in 3D human avatar modeling. Existing methods assume sharp inputs, neglecting the motion blur in animatable avatars and failing to address the detail loss introduced by motion. Our proposed approach integrates a human movement-based motion blur model into 3D Gaussian Splatting. By explicitly modeling human motion trajectories during exposure time, we jointly optimize the trajectories and 3D Gaussians to reconstruct sharp, high-quality human avatars. We employ a pose-dependent fusion mechanism to distinguish moving body regions, optimizing both blurred and sharp areas effectively. Extensive experiments on synthetic and real-world datasets demonstrate that our method significantly outperforms existing methods in rendering quality and quantitative metrics, producing sharp avatar reconstructions and enabling real-time rendering under challenging motion blur conditions.
Index Terms:
All-in-Focus synthesis, main/ultra-wide camera, occlusion-aware networks
I Introduction
@@
Current human avatar modeling approaches neglect motion blur, which can severely degrade the quality of reconstructed avatars by obscuring dynamic human motion details. Several works have proposed to reconstruct static scenes with motion-blurred multi-view inputs, restoring a sharp NeRF or 3DGS by deformable kernels [21, 22, 23] and camera motion model [24, 25]. A dynamic pipeline [26] aims to reconstruct dynamic scenes from blurred monocular videos, but its deblurring paradigm is the same as the previous pipelines [24, 25]. These methods perform well on their datasets, however, these methods, including the dynamic one [26], do not incorporate specialized representations of human motion and would fail if the human movement is more prominent and complex than the camera movement.
To address motion blur in monocular dynamic human videos, we aim to reconstruct sharp 3D human avatars by explicitly modeling human movements. We propose an animatable avatar framework to incorporate a human movement-based motion blur model into 3DGS. We integrate human motion dynamics into 3DGS training by utilizing an interpolation function to model virtual human motion trajectories over the camera’s exposure period. During training, these motion trajectories are optimized jointly with the Gaussians, enabling the reconstruction of sharp human avatars despite motion blur. Specifically, the trajectory of each motion-blurred image is represented by the initial and final human body poses at the beginning and end of the exposure time, respectively. Given that the exposure time is relatively short, we interpolate between these initial and final poses to determine each intermediate pose along the trajectory. Using this trajectory, we generate a sequence of virtual sharp images by deforming the Gaussians with non-rigid and rigid transforms and projecting the Gaussians onto the image plane. These virtual sharp images are then averaged to synthesize the blurred image, adhering to the physical principles of motion blur. Finally, we observe that human movement does not involve the whole body for each frame. Therefore, we propose pose-dependent fusion to effectively distinguish the blurred regions for each pose in a frame, so that our model can focus on modeling motion blur in these blurred areas.
To address the absence of existing human avatar motion deblurring datasets, we collect synthetic and real-world datasets and conduct an extensive analysis to demonstrate that our method effectively addresses motion-blurred inputs and renders high-quality sharp avatars, outperforming existing methods. Additionally, we conduct ablation studies to validate the effectiveness of each component. In summary, our main contributions are as follows.
@@
    We present the first framework designed to recover sharp animatable human avatars from monocular videos with motion blur.
@@ II Related Work
A disadvantage of NeRF-based methods is their inefficiency in training and rendering. 3D Gaussian Splatting [12, 51] (3DGS) shows superiority in quality and efficiency, so 3DGS is integrated into human avatars [52, 13, 14, 15, 16, 17, 18, 19, 53, 54, 55]. KL dispersion [56] and a segmentation plus scaling strategy [57] are used to manipulate point cloud, and kinematic information [58] is used for surface deformation. Non-rigid deformations [20, 59] are integrated into 3DGS to achieve a balance between quality and speed. Lifted optimization and a novel differentiable shading are introduced  [59, 60] to optimize Gaussian parameters while projecting them onto meshes. Contrastive learning [61] and semantic priors [62] are used to capture details and maintain consistency.
@@
Motion Deblurring. Traditional single-image deblurring models the blur with normalized gradient sparsity [63], latent structure priors [64], unified probabilistic model [65], and rotational velocity [66]. Several methods [67, 68, 69, 70] train neural networks on datasets [71], exploring attention mechanisms [72, 73] and multi-scale pipelines [74, 75]. On the other hand, video deblurring harnesses temporal information among different frames to ensure consistency. Conventional approaches employ optical flow to model inter-frame relationships [76, 77]. Several approaches stack neighboring frames as the input of neural networks [78, 79]. With the new proposed benchmark dataset [80], attention mechanisms [81, 82], motion magnitude [83], and wavelet-based [84] methods are proposed to improve performance and efficiency. While these methods yield excellent results, they assume the blur is predominantly from camera movement, and the method performs poorly on blur from human motion. Furthermore, these methods operate in image space and fail to capture scene geometry in $3$D space.
@@
Recently NeRF has demonstrated impressive results in novel view synthesis, and many methods have extended NeRF from sharp static scenes to motion-blurred static [21, 22, 24, 85] and dynamic scenes [26]. However, the static pipelines are not fit for dynamic motion, and the dynamic pipeline [26] is designed for the entire scene, so only the general human motion of the scene can be deblurred. The method lacks specialized human motion representation and would fail if the human movement is more prominent and complex than the camera movement. To improve efficiency, 3DGS is also introduced to model the blur with the covariance of 3D Gaussian [86, 87, 23] and camera motions [88, 25, 89]. Current methods, including the dynamic deblurring pipeline, are designed specifically for motion blur from camera motions instead of human motions, and would not suffice for our aim: restoring from motion blur due to human movements in human avatar modeling.
III Method
@@
Figure 2: Pipeline of our method. For each frame with its corresponding body pose as input, we introduce human motion trajectory modeling to predict human poses $\theta_{start}$ and $\theta_{end}$ at the start $T_{start}$ and end $T_{end}$ of the exposure period of the frame. Then we use interpolation to obtain a virtual human pose sequence. With the pose of each virtual timestamp, we perform non-rigid deformation and rigid transformation to the 3D Gaussian initialized in canonical space. The observation space virtual 3D Gaussians and their colors are accumulated through differentiable rasterization to render sharp virtual images. These images are averaged as Eq. 6 to produce the blurred result. The blurred image is then blended with the virtual sharp result by pose-dependent fusion to render the final output, supervised by the captured input RGB image. During testing, Gaussians can be rendered directly to produce a sharp image without motion trajectory modeling and pose-dependent fusion, enabling fast rendering.
@@
We propose a human avatar modeling pipeline to recover sharp novel views and poses from a motion-blurred monocular video. In the following, we first revisit the concept of 3DGS and the principle of motion blur. In section III-C, we propose human motion trajectory modeling to account for human movements in the exposure period. We show how to perform rigid and non-rigid deformation from the canonical space in section III-D. Then we apply pose-dependent fusion to predict the blurred region of each view, blending the rendered blur region and the rendered sharp region to get optimal results in section III-E.
III-A Preliminary: 3D Gaussian Splatting
@@
III-B Preliminary: Motion Blur Formation
@@
III-C Human Motion Trajectory Modeling
@@
We aim to estimate both $\theta_{\text{start}}$ and $\theta_{\text{end}}$ for each frame, along with the learnable parameters of Gaussians $G$. To predict both poses from a single input, we initialize $\theta_{\text{start}}$ and $\theta_{\text{end}}$ with the input body pose parameters. The learnable poses are used for subsequent rigid transformation, allowing direct optimization through backpropagation. For very small angles (i.e., when $\sin(\phi)$ is close to zero), we assume the corresponding frame has very little motion blur and linear interpolation is used as an approximation. We also predict the sequences of per-frame translations with direct linear interpolation and apply Slerp to the global orientation of the body to get a virtual sequence. We set the virtual frame number $n=5$ as the number of images during an exposure time, balancing training efficiency and rendering quality.
III-D Human Deformation and Color Prediction
@@
III-E Pose-dependent Fusion
@@
where this mask is used to blend the virtual sharp image $\mathbf{C}(x,j)$ with the blurred $\mathbf{B}(x,j)$ (stacked from Eq. 6):
@@
III-F Optimization
We learn the Gaussians and the human motion trajectories for each frame from the loss functions, consisting of an RGB loss $L_{\text{rgb}}$ and perceptual loss $L_{\text{percept}}$ for $\mathbf{C}_{\text{out}}(x,j)$ and $\mathbf{B}_{\text{gt}}$, a mask loss $L_{\text{mask}}$ for foreground human. Skinning weight regularization $L_{\text{skin}}$ and isometric regularization $L_{\text{isopos}},L_{\text{isocov}}$ [20] are implemented to alleviate noisy deformation. The isometric regularization losses regularize both position and covariance $L_{\text{isopos}}$ and $L_{\text{isocov}}$. For the full loss function $L_{\text{full}}=L_{\text{rgb}}+\lambda_{1}L_{\text{percept}}+\lambda_{2}L_{\text{mask}}+\lambda_{3}L_{\text{skin}}\\ +\lambda_{4}L_{\text{isopos}}+\lambda_{5}L_{\text{isocov}}$, $\lambda_{1}=0.01$, $\lambda_{2}=0.1$, $\lambda_{3}$ is set to 10 for initialization and decrease in training. $\lambda_{4}=1$, $\lambda_{5}=100$.
Implementations. Our model is trained for $15k$ iterations on the ZJU-MoCap-Blur dataset and the Real-Human-Blur dataset in $40$ minutes on a single NVIDIA RTX 3090 GPU. We employ the Adam optimizer [92] to optimize our model and the per-frame latent codes. The learning rate of 3D Gaussians is unaltered as the original 3DGS implementation [12]. We initialize the canonical 3D Gaussians with N = $50k$ random samples on the SMPL mesh surface in canonical pose. During optimization, we follow the original 3DGS pipeline [12] to densify and prune the 3D Gaussians, using the view-space positional gradients to determine the densification regions. To alleviate overfitting in color prediction, inverse rigid transformation is implemented in the same way as  [20] to canonicalize the viewing direction. Then the spherical harmonics basis of canonicalized viewing direction and pose-dependent feature from non-rigid MLP are used as inputs for the color MLP.
We begin to learn the human trajectory modeling at $3k$ iterations, then after $7k$ iterations, we start learning the mask for training stabilization. Before $7k$ iterations, we use the averaged blurred rendered result as input of the loss function. $f_{\theta_{\text{nr}}}$ is an MLP with $3$ hidden layers of $128$ dimensions, and $f_{\theta_{\text{rigid}}}$ is an MLP with $4$ hidden layers of $128$ dimensions as previous works. The color MLP consists of a single $64$-dimensional hidden layer, which has been proven effective [20]. The mask prediction network $f_{\theta_{\text{fuse}}}$ is an MLP with $4$ hidden layers of $64$ dimensions.
IV Experiments
@@
Figure 3: Qualitative results on ZJU-MoCap-Blur. Compared with existing human avatar methods, our method generates sharper novel views and preserves more details with original motion-blurred inputs.
Figure 4: Qualitative results on Real-Human-Blur dataset. Our method generates sharper novel poses that are more faithful and preserve more details.
Figure 5: Qualitative results on ZJU-MoCap-Blur (pre-deblurred). The input sequences are deblurred by state-of-the-art video deblurring method [84]. Compared with existing human avatar methods, our method generates sharper novel views and preserves more details with original motion-blurred inputs.
Table I: Quantitative results on ZJU-MoCap-Blur. The input sequence is the motion-blurred images. The best performance is in boldface, and the second best is underlined. We present the training time as GPU and the rendering speed as FPS.
@@
Table II: Quantitative results on ZJU-MoCap-Blur (pre-deblurred). The input sequences are pre-deblurred by state-of-the-art video deblurring method [84]. The best performance is in boldface, and the second best is underlined. We present the training time as GPU and the rendering speed as FPS.
@@ IV-A Datasets
ZJU-MoCap-Blur. This is the main dataset for quantitative evaluation. We pick six sequences (377, 386, 387, 392, 393, 394) from the ZJU-MoCap dataset and follow the training/test split of HumanNeRF [42]. We synthesize motion blur as the pipeline of current benchmark datasets on motion deblurring [67, 93, 94]. ZJU-MoCap is shot at a frame rate of 60 fps. To synthesize realistic motion blur without artifacts like previous work [67], we increase the video frame rate to 480 fps using a state-of-the-art frame interpolation method [95]. Then we average this sharp high frame rate successive frames to generate a blurry image to approximate a long exposure time, and the generated frames are temporally centered on a real-captured ground truth frame from original ZJU-MoCap. We apply varying blur sizes for the dataset, so the dataset consists of scenes with small blur(17 frames to synthesize one frame) for sequences 377 and 392, medium blur (33 frames) for 393 and 394, and large blur (49 frames) for sequences 386 and 387. To make the dataset more realistic, we use EasyMocap [96, 47] to re-calculate the human poses and the human masks of the synthesized blurred image sequences. We train and evaluate the Real-Human-Blur dataset at the resolution of $540\times 540$, $960\times 540$, and $360\times 640$, depending on the original resolution of the captured video. Following previous works [20, 42], we conduct quantitative evaluations on novel view synthesis and show qualitative results for animation on out-of-distribution poses. LPIPS in all the tables are scaled up by 1000.
@@
Real-Human-Blur. Given the absence of publicly available, real-world datasets that are specifically tailored for tackling the challenge of motion blur in human avatar modeling, we have curated a dataset consisting of monocular motion-blurred videos. The videos were recorded using a high-resolution DSLR camera under varying lighting and background settings, allowing for rich visual details and realistic blur patterns caused by human motion. We use SPIN [97] to obtain approximate body poses and employ SAM [98] for segmenting the foreground human. Since there is no accurate sharp ground truth for these real captures, we use this dataset solely for qualitative comparisons on novel pose synthesis.
IV-B Baseline Comparisons